OpenStudy (anonymous):

n^2-12=4n I don't know how to solve for n!

OpenStudy (anonymous):

Subtract 4n from both sides of the equation so you have n^2 - 4n - 12 = 0. Factor.

OpenStudy (anonymous):

\[(n-6)(n+2)=0\]So either n-6=0 or n+2=0 will make this expression true. Solving n-6=0, n=6; solving n+2=0, n=-2. The solutions are n=6, n=-2

OpenStudy (anonymous):

For quadratic of form x^2 + bx + c , u can try to find 2 numbers p and q that add up to b (-4 in this case) and multiply together to give c (-12 in this case). Then the factors are (x+p)(x+q). Often works with made up problems, else use quadratic formula.